# The Impact of Inference Acceleration Strategies on Bias of LLMs

### v1

@@ Abstract
Last few years have seen unprecedented advances in capabilities of Large Language Models (LLMs). These advancements promise to deeply benefit a vast array of application domains. However, due to their immense size, performing inference with LLMs is both costly and slow. Consequently, a plethora of recent work has proposed strategies to enhance inference efficiency, e.g., quantization, pruning, and caching. These acceleration strategies reduce the inference cost and latency, often by several factors, while maintaining much of the predictive performance measured via common benchmarks. In this work, we explore another critical aspect of LLM performance: demographic bias in model generations due to inference acceleration optimizations. Using a wide range of metrics, we probe bias in model outputs from a number of angles. Analysis of outputs before and after inference acceleration shows significant change in bias. Worryingly, these bias effects are complex and unpredictable. A combination of an acceleration strategy and bias type may show little bias change in one model but may lead to a large effect in another. Our results highlight a need for in-depth and case-by-case evaluation of model bias after it has been modified to accelerate inference.
@@ 1 Introduction
Modern-day LLMs like LLaMA and GPT-4 show remarkable language generation capabilities, leading to a surge in their popularity and adoption (Ziems et al., 2024; Wei et al., 2022; Bubeck et al., 2023). However, owing to their immense size, deploying these models can be challenging, or even infeasible for consumer-grade devices. A flurry of research has proposed acceleration strategies such as quantization and pruning to enable efficient inference (Park et al., 2024; Zhu et al., 2023). The objective of these strategies is typically to reduce the model size while maintaining predictive performance. Over time, these strategies have become increasingly prevalent with integration into popular libraries like HuggingFace (Hug, 2024) and libraries such as vLLM (Kwon et al., 2023).
@@
This paper explores how the demographic bias in the model output changes after the implementation of inference acceleration strategies. Specifically, we aim to answer the following research questions:
@@
Table 1: [Mistral-7B model on DiscrimEvalGen data] Decisions by the base model and its 4-bit weight quantized version. The choice and reasoning changes from the base to the quantized version.
@@ 2 Related Work
The effect of model size on trust criteria. Recent work has started exploring the impact of model size on trust related criteria. For example, Perez et al. (2022) find that larger models tend to overly agree with user views. Sun et al. (2024) show that smaller models can reduce privacy risks. Huang et al. (2024) find that smaller models are more vulnerable to backdoor attacks. Mo et al. (2024) find that larger models are more susceptible to manipulation through malicious demonstrations. Jaiswal et al. (2024) offer a fine-grained benchmark for evaluating the performance of compressed LLMs on more intricate, knowledge-intensive tasks such as reasoning, summarization, and in-context retrieval. By measuring perplexity, they show that pruned models suffer from performance degradation, whereas quantized models tend to perform better. Xu and Hu (2022) find that knowledge distillation causes a monotonic reduction in toxicity in GPT-2, though it shows only small improvements in reducing bias on counterfactual embedding-based datasets. These analyses differ from the current paper in one of the following two ways: (i) they are limited to less recent, pre-trained models, which may not adequately represent the complexities of modern LLMs with significantly more parameters; (ii) they target trustworthiness desiderata beyond bias, e.g., backdoor attacks.
Effect of inference acceleration strategies on trustworthiness. Gonçalves and Strubell (2023) measure the impact of quantization and knowledge distillation on LLMs, and show that longer pretraining and larger models correlate with higher demographic bias, while quantization appears to have a regularizing effect. The bias metrics they consider focus on embeddings or token output probabilities, while we consider a larger range of metrics that focus on properties of generated texts. Hong et al. (2024), in a follow-up to Wang et al. (2024), provide a broader assessment of trustworthiness under compression strategies like quantization and pruning, including adversarial settings. However, their study relies on a single metric to evaluate stereotype bias, which may not capture the broader complexity of bias. We, on the other hand, aim to provide a comprehensive evaluation of bias across multiple dimensions to better understand the impact of inference acceleration strategies. Finally, while these previous benchmarks show largely uniform and predictable effects of inference acceleration on bias, by leveraging a richer set of metrics, our analysis shows a much more nuanced picture and a need for case-by-case evaluation.
@@ 3 Measuring Bias in LLM Outputs
We leave out embedding-based metrics from our analysis since (i) the more typical use-case of modern, instruction-tuned, LLMs like LLaMA and GPT-4 is prompt-tuning or fine-tuning rather than adapting the models using embeddings and (ii) embedding bias is not guaranteed to lead to bias in the text generations. We initially considered classification-based bias metrics (e.g.Dhamala et al.), which assess differences in measures like toxicity and sentiment on common datasets like Wikipedia. Preliminary analysis showed very little overall toxicity in model outputs, most likely due to heavy alignment on these datasets. For this reason, we did not further consider these metrics.
@@
CrowSPairs (Nangia et al., 2020) is a dataset of crowd-sourced sentence pairs designed to evaluate stereotypes related to race, gender, sexual orientation, religion, age, nationality, disability, physical appearance, and socioeconomic status. Each pair consists of one sentence that demonstrates a stereotype and the other that demonstrates the opposite of the stereotype. Given a pair $(s_{\text{more}},s_{\text{less}})$ where $s_{\text{more}}$ is presumed to be more stereotypical, the metric measures $\mathbb{I}[p(s_{\text{more}})>p(s_{\text{less}})]$ and averages this quantity over all pairs, with $\mathbb{I}$ as the indicator function. The score ranges from $[0,1]$.
GlobalOpinionQA (Durmus et al., 2024) uses multiple-choice questions to assess the opinions stated by a model relative to aggregated population opinions from different countries. The goal is to identify biases the model may have in representing diverse viewpoints. We follow the measurement procedure of Durmus et al. with one exception: we use the Wasserstein Distance as our similarity metric (leveraging the implementation provided by the Python scipy library (Virtanen et al., 2020)). Durmus et al. use 1-Jensen-Shannon Distance as a similarity metric, which can become highly skewed when the distributions have very little or no overlap. In contrast, Wasserstein Distance is more sensitive to the geometry of the probability distributions Arjovsky et al. (2017). The bias value is then the Gini coefficient of the Wasserstein Distance for each country. The metric lies in range $[0,1]$. The dataset does not provide responses from all countries to all questions, making it difficult to analyze overall value tendencies consistently. To address this, we exclude any countries that do not have responses to at least 50 questions from our analysis.
@@
DiscrimEval (Tamkin et al., 2023) consists of 70 hypothetical decision making scenarios, e.g., approving a loan. For each scenario, the model is prompted to make a binary yes/no decision about a person described in terms of age, gender and race (9, 3 and 5 choices, respectively). A yes decision is always advantageous. Following Tamkin et al., we append "My answer would be" to the prompt to steer the generations towards producing binary decisions and record the model’s softmax probability of “yes” or “no” being generated as the first token. For a scenario $q_{i}$ and a set of demographic groups $G$ ($9\times 3\times 5=135$ intersectional groups in this case), we repeatedly reformulate $q_{i}$, substituting the demographic information for all groups $g\in G$ one by one, and measure the difference between the highest and lowest probability of “yes” for all groups $g\in G$. Specifically, the bias score is:
@@
where $Q$ is the set of all questions and $n=|Q|$. We use the dataset’s “explicit” version in our evaluation, directly including demographic attributes in the prompt rather than implying it via names.
DiscrimEvalGen. The original design of DiscrimEval evaluates bias by analyzing the probability of the first token being “yes” or “no”, reducing the model’s output to a simplified binary decision. However, this approach (i) considers only a single token for bias measurement ignoring the subsequent tokens and (ii) overlooks the model’s broader preferences among demographic groups. With the aim of overcoming these issues, we propose a new dataset DiscrimEvalGen. Whereas DiscrimEval asks the same question separately for each demographic group $g$, DiscrimEvalGen forces the model to make a single choice. Specifically, we (i) present the question to the model and describe that the candidates are persons from different groups, e.g., a man, a woman, a non-binary person; (ii) describe that the benefit (e.g., a work contract) can be granted to only a single person; and (iii) ask the model to make its choice. Let $q\in Q$ be the questions, $g\in G$ be the groups, and $n_{g}$ be the number of times a group is selected by the model with $\sum_{g\in G}n_{g}=|Q|$, then the bias metric is:
@@ 4 Experimental Setup
Models and Infrastructure. We analyze three different models: LLaMA-2 (Touvron et al., 2023), LLaMA-3.1 (Dubey et al., 2024), and Mistral-0.3 (Jiang et al., 2023). We consider the smallest size variant of each model: LLaMA-2-7B, LLaMA-3.1-8B, and Mistral-7B-v0.3 (license information in Section A). These models were selected due to their recency, widespread use, and compatibility with our resource constraints, which included a single node equipped with four NVIDIA A100 GPUs that was shared among several research teams. Our evaluation focuses on the chat versions of these models, which are specifically designed to align with human values and preferences. We used the GitHub Copilot IDE plugin to assist with coding.
@@
    Activation-aware Weight Quantization (AWQ) (Lin et al., 2024) quantizes the parameters by taking into account the data distribution in the activations produced by the model during inference. We use the 4-bit version and the authors do not provide a 8-bit implementation.
@@
Pruning removes a subset of model weights to reduce the high computational cost of LLMs while aiming to preserve performance. Traditional pruning methods require retraining (Cheng et al., 2024). More recent approaches prune weights post-training in iterative weight-update processes, e.g., SparseGPT (Frantar and Alistarh, 2023). We use the Wanda method by Sun et al. (2024) which uses a pruning metric based on both weight magnitudes and input activation norms. The sparse model obtained after pruning is directly usable without further fine-tuning. We consider two variants: (i) Unstructured Pruning (WU) with a $50\%$ sparsity ratio, removing half of the weights connected to each output; and (ii) Structured Pruning (WS), enforcing structured N:M sparsity where at most N out of every M contiguous weights are allowed to be non-zero, allowing the computation to leverage matrix-based GPU optimizations. We use a $2:4$ compression rate. Prior work has shown that pruned models can maintain comparable performance even at high compression rates (Sun et al., 2024; Jaiswal et al., 2024; Frantar and Alistarh, 2023), including the $2:4$ rate used here.
Parameters. As described in Section 3, most bias metrics are designed such that they only support greedy decoding, resulting in deterministic outputs. Only DT-Stereotyping and DiscrimEvalGen support both stochastic decoding and greedy decoding. With stochastic decoding, we sample the output $5$ times and report the average bias. The models can be used with and without the developer-prescribed instruction templates (using special tokens for instruction boundaries). While instruction formats can have an unpredictable impact on the model performance (Fourrier et al., 2023), instruction templates’ impacts on model bias are less well understood. We thus study both configurations, with and without the instruction template. The main paper includes the results without the template, while results with instructions templates are shown in Appendix C.
@@
LLaMA-2 0.6 $\uparrow$3 0.62 $\downarrow$2 0.59 0.6 0.6 0.6 $\downarrow$2 0.59 $\downarrow$2 0.59
Mistral 0.55 $\uparrow$13 0.62 $\uparrow$9 0.6 $\uparrow$2 0.56 0.55 $\downarrow$4 0.53 ni ni
LLaMA-3.1 0.58 $\downarrow$9 0.53 $\uparrow$12 0.65 $\downarrow$3 0.56 $\uparrow$2 0.59 $\downarrow$2 0.57 $\downarrow$2 0.57 $\downarrow$2 0.57
@@ 5 Results
Table 2 shows the bias of base models w.r.t. each metric, and the change in bias as a result of inference acceleration. We show examples of generations and further output characteristics in the Appendix. The table shows that inference acceleration strategies can have significant, albeit nuanced, impacts on bias in LLMs. While some strategies consistently reduce certain biases, others yield mixed results depending on the model and context. The results also show that while the input probability-based metric, CrowSPairs, does not show much change in bias across the board, considering a wider range of metrics paints a much more diverse picture. While the magnitude of changes varies, we largely see similar trends of unpredictable effects on downstream bias both with and without the instruction template (Appendix C). Although we did not track the exact runtime, our experiments took several GPU days. We now analyze each RQ from Section 1 in detail.
@@
Inference acceleration strategies can have disparate effects across different types of bias metrics. Specifically, we observe:
No significant impact on log-likelihood of stereotypical sentences as measured by the CrowsPairs dataset. Most acceleration strategies show little to no significant effect on the log-likelihood of counterfactual sentences. The results are largely in line with Gonçalves and Strubell (2023) who also show a relatively mild effect of quantization on bias measured via CrowSPairs, although they consider the previous generation of LLMs like BERT and RoBERTa. We provide a detailed breakdown of results per bias type in Table B.1.
Minimal impact on values and opinions in the GlobalOpinionQA task. We observe little effect of inference strategies on the values and opinions represented by the models (see 2(b)). Notably, KV cache quantization showed no negative impact at all. While the overall similarity of responses per country remains unaffected, there are still subtle shifts in the ranking of individual countries, as reflected in the world maps in Figure B.1.
Nuanced effects on models’ bias in recalling facts from different countries. In the WorldBench dataset, KV cache quantization showed slight improvements in mean average error, while pruning strategies produced non-uniform effects across different models. We report detailed disparity scores across income groups and regions in Table B.5. Overall, except for pruning, inference acceleration strategies had no notable effect on the models’ factual recall abilities for different countries.
@@
Varying bias patterns in allocation-based decision-making scenarios. In the original DiscrimEval benchmark, structured pruning consistently achieved the lowest bias score across models, followed closely by KV cache quantization. On the other hand, AWQ quantization resulted in a notable increase in bias. We move on to analyze bias in relative decision scenarios for longer text generations. In the DiscrimEvalGen dataset, we observe more significant shifts in resource allocation based on gender, with AWQ leading to increased discrimination across models and sampling strategies. A detailed breakdown of decisions per model and tested attributes in Table B.4 shows that inference acceleration strategies influence the models’ tendency to give no response or refuse an answer. Both Mistral and LLaMA-3.1 display a tendency to favor the non-binary person, though this effect is reduced when pruning strategies are applied.
@@
KV cache quantization and structured Wanda pruning showed promising trends across datasets and models, frequently showing minimal changes or slight improvements in bias scores. However, structured pruning exhibited certain drawbacks. When examining parse rates and no-response rates, we found that this strategy can cause the model to fail to perform the task, follow instructions, or produce nonsensical, repetitive outputs. Overall, our results suggest quantizing weights can have more drastic, unpredictable impacts on bias as compared to KV cache quantization.
@@
The effects of inference acceleration strategies on stereotype agreeability vary markedly across models. A detailed breakdown of agreement, disagreement, and no-response rates for nucleus sampling in Table B.2 illustrates how the models’ baselines already differ. LLaMA models most frequently provide no response, while Mistral shows a higher rate of both agreement and disagreement. The impact of inference acceleration strategies is notably more pronounced for Mistral, with agreements increasing by over $75\%$ relative to the base model for both AWQ and unstructured pruning.

### v2

# The Impact of Inference Acceleration on Bias of LLMs

@@ Abstract
Last few years have seen unprecedented advances in capabilities of Large Language Models (LLMs). These advancements promise to benefit a vast array of application domains. However, due to their immense size, performing inference with LLMs is both costly and slow. Consequently, a plethora of recent work has proposed strategies to enhance inference efficiency, e.g., quantization, pruning, and caching. These acceleration strategies reduce the inference cost and latency, often by several factors, while maintaining much of the predictive performance measured via common benchmarks. In this work, we explore another critical aspect of LLM performance: demographic bias in model generations due to inference acceleration optimizations. Using a wide range of metrics, we probe bias in model outputs from a number of angles. Analysis of outputs before and after inference acceleration shows significant change in bias. Worryingly, these bias effects are complex and unpredictable. A combination of an acceleration strategy and bias type may show little bias change in one model but may lead to a large effect in another. Our results highlight a need for in-depth and case-by-case evaluation of model bias after it has been modified to accelerate inference.
@@
The Impact of Inference Acceleration on Bias of LLMs
Elisabeth Kirsten1,3, Ivan Habernal1,3, Vadant Nanda2, Muhammad Bilal Zafar1,3 1Ruhr University Bochum, 2Aleph Alpha, 3UAR Research Center for Trustworthy Data Science and Security Correspondence: elisabeth.kirsten@rub.de
@@ 1 Introduction
Modern-day LLMs like LLaMA and GPT-4 show remarkable language generation capabilities, leading to a surge in their popularity and adoption (Ziems et al., 2024; Wei et al., 2022; Bubeck et al., 2023). However, owing to their immense size, deploying these models can be challenging, or even infeasible for consumer-grade devices. A flurry of research has proposed acceleration strategies such as quantization and pruning to enable efficient inference (Park et al., 2024; Zhu et al., 2023). The objective of these strategies is typically to reduce the model size while maintaining predictive performance. Over time, these strategies have become increasingly prevalent with integration into popular libraries like HuggingFace (Hug, 2024) and vLLM (Kwon et al., 2023).
@@
In this paper, we explore how demographic bias in the model output changes after the implementation of inference acceleration strategies. Specifically, we aim to answer the following questions:
@@
Table 1: [Mistral-7B model on DiscrimEvalGen data] Decisions made by the base model and its 4-bit weight quantized version. The choice and reasoning changes from the base to the quantized version.
@@
The code for our experiments is available at https://github.com/aisoc-lab/inference-acceleration-bias.
@@ 2 Related Work
The effect of model size on trust criteria. Recent work has started exploring the impact of model size on trust related criteria. For example, Perez et al. (2022) find that larger models tend to overly agree with user views. Sun et al. (2024) show that smaller models can reduce privacy risks. Huang et al. (2024) find that smaller models are more vulnerable to backdoor attacks. Mo et al. (2024) find that larger models are more susceptible to manipulation through malicious demonstrations. Jaiswal et al. (2024) offer a fine-grained benchmark for evaluating the performance of compressed LLMs on more intricate, knowledge-intensive tasks such as reasoning, summarization, and in-context retrieval. By measuring perplexity, they show that pruned models suffer from performance degradation, whereas quantized models tend to perform better. Xu and Hu (2022) find that knowledge distillation causes a monotonic reduction in toxicity in GPT-2, though it shows only small improvements in reducing bias on counterfactual embedding-based datasets. These analyses differ from our paper in one of the following ways: (i) they are limited to less recent, pre-trained models, which may not adequately represent the complexities of modern LLMs with significantly more parameters; (ii) they target trustworthiness desiderata beyond bias, e.g., backdoor attacks.
Effect of inference acceleration on trustworthiness. Gonçalves and Strubell (2023) measure the impact of quantization and knowledge distillation on LLMs, and show that longer pretraining and larger models correlate with higher demographic bias, while quantization appears to have a regularizing effect. The bias metrics they consider focus on embeddings or token output probabilities, while we consider a larger range of metrics that focus on properties of generated texts. Hong et al. (2024), in a follow-up to Wang et al. (2024), provide a broader assessment of trustworthiness under compression strategies like quantization and pruning, including adversarial settings. However, their study relies on a single metric to evaluate stereotype bias, which may not capture the broader complexity of bias. We, on the other hand, aim to provide a comprehensive evaluation of bias across multiple dimensions to better understand the impact of inference acceleration strategies. Finally, while these previous benchmarks show largely uniform and predictable effects of inference acceleration on bias, by leveraging a richer set of metrics, our analysis shows a much more nuanced picture and a need for case-by-case evaluation.
@@ 3 Measuring Bias in LLM Outputs
We leave out embedding-based metrics from our analysis since (i) the more typical use-case of modern, instruction-tuned LLMs like LLaMA and GPT-4 is prompt-tuning or fine-tuning rather than adapting the models using embeddings and (ii) embedding bias is not guaranteed to lead to bias in the text generations. While we initially considered classification-based bias metrics (e.g., those in Dhamala et al.) which consider difference in measures like toxicity and sentiment on common datasets like Wikipedia. A preliminary analysis showed very little overall toxicity in model outputs, most likely due to heavy alignment on these datasets. For this reason, we did not further consider these metrics.
@@
CrowSPairs (Nangia et al., 2020) is a dataset of crowd-sourced sentence pairs designed to evaluate stereotypes related to race, gender, sexual orientation, religion, age, nationality, disability, physical appearance, and socioeconomic status. Each pair consists of one sentence that demonstrates a stereotype and the other that demonstrates the opposite of the stereotype. Given a pair $(s_{\text{more}},s_{\text{less}})$ where $s_{\text{more}}$ is presumed to be more stereotypical, the metric measures $\mathbb{I}[p(s_{\text{more}})>p(s_{\text{less}})]$ and averages this quantity over all pairs. $\mathbb{I}$ denotes the indicator function. The resulting score is in the range $[0,1]$.
GlobalOpinionQA (Durmus et al., 2024) uses multiple-choice questions to assess the opinions stated by a model relative to aggregated population opinions from different countries. The goal is to identify model bias in representing diverse viewpoints. We follow the same measurement procedure as Durmus et al. with one exception: we use the Wasserstein distance as our divergence metric. Durmus et al. use 1-Jensen-Shannon distance, which can become highly skewed when the distributions have very little or no overlap. In contrast, the Wasserstein distance is more sensitive to the geometry of the probability distributions Arjovsky et al. (2017). We compute the Wasserstein distance using the implementation provided by the Python scipy library (Virtanen et al., 2020).
The final bias value is then the Gini coefficient of the Wasserstein distance for each country. The metric lies in the range $[0,1]$. The dataset does not provide responses from all countries to all questions, making it difficult to analyze overall value tendencies consistently. To address this, we exclude countries that do not have responses to at least 50 questions from our analysis.
@@
DT–Stereotyping. DecodingTrust (Wang et al., 2024) is a framework for evaluating the trustworthiness of LLMs across eight dimensions, one of which is stereotype bias. The dataset consists of custom-made statements (from now on referred to as DT-Stereotyping) designed to provoke unsafe (e.g., biased, toxic) responses. Following Wang et al. (2024), we ask the model to either agree or disagree with these statements. To measure stereotyping behavior, we compute the fraction of times the model agrees with the stereotypical statements $\frac{n_{\text{agree}}}{n}\in[0,1]$. The original dataset is evaluated in three evaluation scenarios that instruct the model to behave as a helpful assistant (benign setting), in disregard of its content policy (untargeted), and with targeted bias towards the target group (targeted). We use the untargeted setting to evaluate the resilience of model alignment, without explicitly targeting any group adversarially.
@@
DiscrimEval (Tamkin et al., 2023) consists of 70 hypothetical decision making scenarios, e.g., approving a loan. For each scenario, the model is prompted to make a binary yes/no decision about a person described in terms of age, gender and race (9, 3, and 5 choices, respectively). A “yes” decision is always advantageous. Following Tamkin et al., we append "My answer would be" to the prompt to steer the generations towards producing binary decisions and record the model’s softmax probability of “yes” or “no” being generated as the first token. For a scenario $q_{i}$, and a set of demographic groups $G$ ($9\times 3\times 5=135$ intersectional groups in this case), we repeatedly reformulate $q_{i}$, substituting the demographic information for all groups $g\in G$ one by one, and measure the difference between the highest and lowest probability of “yes” for all groups $g\in G$. Specifically, the bias score is:
@@
where $Q$ is the set of all questions and $n=|Q|$. We use the “explicit” version of the dataset in our evaluation, directly including demographic attributes in the prompt rather than implying it via names.
DiscrimEvalGen. The original design of DiscrimEval evaluates bias by analyzing the probability of the first token being “yes” or “no”, reducing the model’s output to a simplified binary decision. This approach has the following limitations:
  1. 1.
    It only considers the first output token, ignoring the rest of the generation.
  2. 2.
    Even at the first token, the bias is computed by considering the Softmax probabilities of ‘yes’ and ‘no’ which could be miscalibrated and may not adequately capture the model’s uncertainty  (Cruz et al., 2024).
  3. 3.
    The model is required to independently evaluate each person and could potentially assign advantageous outcomes to everyone regardless of their demographic features. Such an unconstrained setup may not test more subtle inclinations of the model, e.g., preferences when required to make a choice between different demographic groups or refusing to make a choice at all. In fact, Bai et al. (2024) show that GPT-4 assigns benefits to various demographic groups in similar proportions.
With the aim of overcoming these issues, we propose a new dataset DiscrimEvalGen. While DiscrimEval asks the same question separately for each demographic group $g$, DiscrimEvalGen forces the model to make a single choice. Specifically, we (i) present the question to the model and describe that the candidates are persons from different groups, e.g., a man, a woman, a non-binary person; (ii) describe that the benefit (e.g., a work contract) can be granted to only a single person; and (iii) ask the model to make its choice.
Let $q\in Q$ be the questions, $g\in G$ be the groups, and $n_{g}$ be the number of times a group is selected by the model with $\sum_{g\in G}n_{g}=|Q|$, then the bias metric is:
@@ 4 Experimental Setup
Models and Infrastructure. We analyze three different models: LLaMA-2 (Touvron et al., 2023), LLaMA-3.1 (Dubey et al., 2024), and Mistral-0.3 (Jiang et al., 2023). We consider the smallest size variant of each model: LLaMA-2-7B, LLaMA-3.1-8B, and Mistral-7B-v0.3 (license information in Appendix A). These models were selected due to their recency, widespread use, and compatibility with our resource constraints, which included a single node equipped with four NVIDIA A100 GPUs that was shared among several research teams. Our evaluation focuses on the chat versions of these models, which are specifically designed to align with human values and preferences. We used the GitHub Copilot IDE plugin to assist with coding.
@@
    Activation-aware Weight Quantization (AWQ) (Lin et al., 2024) quantizes the parameters by taking into account the data distribution in the activations produced by the model during inference. We use the 4-bit version as the authors do not provide an 8-bit implementation.
@@
Pruning removes a subset of model weights to reduce the high computational cost of LLMs while aiming to preserve performance. Traditional pruning methods require retraining (Cheng et al., 2024). More recent approaches prune weights post-training in iterative weight-update processes, e.g., SparseGPT (Frantar and Alistarh, 2023). We use the Wanda method by Sun et al. (2024) which uses a pruning metric based on both weight magnitudes and input activation norms. The sparse model obtained after pruning is directly usable without further fine-tuning. We consider two variants: (i) Unstructured Pruning (WU) with a sparsity ratio of $50\%$, eliminating $50\%$ of the weights connected to each output; and (ii) Structured Pruning (WS) which induces a structured N:M sparsity, where at most N out of every M contiguous weights are allowed to be non-zero, allowing the computation to leverage matrix-based GPU optimizations. We use a $2:4$ compression rate. Prior work has shown that pruned models can preserve comparable performance levels even at high compression rates (Sun et al., 2024; Jaiswal et al., 2024; Frantar and Alistarh, 2023), e.g., $2:4$ considered here.
Parameters. As described in Section 3, most bias metrics are designed such that they only support greedy decoding, resulting in deterministic outputs. Only DT-Stereotyping and DiscrimEvalGen support stochastic decoding in addition to greedy decoding. When using stochastic decoding, we sample the output $5$ times and report the average bias.
The models can be used with and without the developer-prescribed instruction templates (with special tokens specifying the start and end of instructions). Past investigations have shown that instruction and answer formats can have an unpredictable impact on the model performance (Fourrier et al., 2023; Alzahrani et al., 2024). However, the impact of using or not using the instruction template on model bias is less well understood. We thus study both configurations. The main paper includes the results without the template, while results with instructions templates are shown in Appendix C.
@@
LLaMA-2 0.52 $\uparrow$8 0.56 $\uparrow$6 0.55 0.52 $\uparrow$2 0.53 $\downarrow$8 0.48 $\downarrow$10 0.47 $\downarrow$12 0.46
Mistral 0.43 $\uparrow$37 0.59 $\uparrow$19 0.51 $\uparrow$2 0.44 0.43 $\downarrow$21 0.34 ni ni
LLaMA-3.1 0.4 $\uparrow$25 0.5 $\uparrow$38 0.55 $\uparrow$10 0.44 $\uparrow$5 0.42 $\downarrow$3 0.39 0.4 $\downarrow$3 0.39
@@ 5 Results
Table 2 shows the bias of base models w.r.t. each metric, and the change in bias as a result of inference acceleration. We show examples of generations and further output characteristics in the Appendix. The table shows that inference acceleration strategies can have significant, albeit nuanced, impacts on bias in LLMs. While some strategies consistently reduce certain biases, others yield mixed results depending on the model and context. The results also show that while the input probability-based metric, CrowSPairs, does not show much change in bias across the board, considering a wider range of metrics paints a much more diverse picture. While the exact magnitude of changes varies, we largely see similar trends of unpredictable effects on downstream bias both with and without the instruction template (Appendix C). Although we did not track the exact runtime, our experiments required several GPU days to complete. We now analyze each RQ from Section 1 in detail.
@@
Inference acceleration strategies have disparate impacts on different types of bias metrics. Specifically, we see:
No significant impact on log-likelihood of stereotypical sentences as measured by the CrowsPairs dataset. Most acceleration strategies show little to no significant effect on the log-likelihood of counterfactual sentences. The results are largely in line with Gonçalves and Strubell (2023) who also show a relatively mild effect of quantization on bias measured via CrowSPairs, although they consider the previous generation of LMs like BERT and RoBERTa. We provide a detailed breakdown of results per bias type in Table B.1. Structured pruning leads to small improvements in bias scores for certain bias types, such as "age". However, the average improvements over the base model across all bias types are modest, generally less than 10%.
Subtle shifts in values and opinions in the GlobalOpinionQA task. We observe little effect of inference strategies on the values and opinions represented by the models (see 2(b)). AWQ quantization increased bias across all models, with changes of up to 36%. Structured pruning also leads to noticeable shifts, including a 45% increase in bias scores for Mistral. Despite these changes, overall bias scores remain low, suggesting that the general similarity of responses across countries is largely unaffected. Notably, KV cache quantization shows no negative impact. While the overall similarity of responses per country often remains stable, there are still subtle shifts in the ranking of individual countries, as reflected in the world maps in Figure B.1.
Pruning influences models’ ability to recall country-specific facts. In the WorldBench dataset, 8-bit and KV cache quantization showed improvements in mean average error, whereas pruning strategies and AWQ quantization increased bias scores. We report detailed disparity scores across income groups and regions in Table B.5. Pruning leads to higher disparities across regions and income groups in 8/12 cases. In contrast, other inference acceleration methods had non-uniform or minimal influence on models’ factual recall performance across countries.
@@
Varying bias patterns in allocation-based decision-making scenarios. In DiscrimEval, structured pruning consistently achieved the lowest bias score across models, followed closely by KV cache quantization. On the other hand, AWQ quantization resulted in a notable increase in bias.
In the DiscrimEvalGen dataset, which measures bias in relative decision making scenarios and longer text generations, we observe more significant shifts in resource allocation based on gender, with AWQ leading to increased bias across models and sampling strategies. A detailed breakdown of decisions per model and tested attributes in Table B.4 shows that inference acceleration strategies influence the models’ tendency to give no response or refuse an answer. Both Mistral and LLaMA-3.1 display a tendency to favor the non-binary person, though this effect is reduced when pruning strategies are applied.
@@
KV cache quantization and structured Wanda pruning showed promising trends across datasets and models, frequently showing minimal changes or slight improvements in bias scores. However, structured pruning exhibited certain drawbacks. When examining parse rates and no-response rates, we found that this strategy can cause the model to fail to perform the task, follow instructions, or produce nonsensical, repetitive outputs. Overall, our results suggest quantizing weights can have more drastic, unpredictable impacts on bias compared to KV cache quantization.
@@
The effects of inference acceleration strategies on stereotype agreeability vary markedly across models. A detailed breakdown of agreement, disagreement, and no-response rates for nucleus sampling in Table B.2 illustrates how the models’ baselines already differ. LLaMA models most frequently provide no response, while Mistral shows a higher rate of both agreement and disagreement. Notably, the impact of inference acceleration strategies is much more pronounced for Mistral, with agreements increasing by over $75\%$ relative to the base model for both AWQ and unstructured pruning.